OpenStudy (anonymous):

f(x) = square root ( x^4 -16x^2) let f be the function given by: a) Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5

OpenStudy (rulnick):

f(x) = sqrt ( x^2 (x+4) (x-4) ) domain is all real numbers >= 4

OpenStudy (rulnick):

(at least, that's the domain if you are not yet doing complex numbers -- are you doing complex numbers yet?)

OpenStudy (anonymous):

i dont need to find complex numbers for this

OpenStudy (anonymous):

i got -4 to be a domain too

OpenStudy (rulnick):

No, what I mean is - if you are considering only real-valued functions, then the domain is all real numbers >= 4 (I guess this is the case) - you will learn later that for complex-valued functions, the domain becomes all real numbers (you're probably not there yet).

OpenStudy (rulnick):

So I think the answer you want is the one I gave at the top.

OpenStudy (anonymous):

how do you do part c ?

OpenStudy (campbell_st):

If you are using the set of real numbers then the domain is \[x \le -4, x \ge 4\] for the derivative \[f'(x) = 1/2(x^4 - 16x^2)^{-1/2} \times (4x^3 - 32x)\] find the gradient of the tangent (m) at x =5 by substituting x = 5. Then the the gradient of the normal is -1/m

OpenStudy (rulnick):

Once you get (b), which is a small mess but doable, just evaluate f'(5) and take the negative reciprocal.

OpenStudy (campbell_st):

with the domain.. be careful and (-4)^ - 16(-4)^2 gives the same answer as (4)^4 - 16(4)^2

OpenStudy (campbell_st):

so the domain is \[x \le -4, x \ge 4\]

OpenStudy (anonymous):

f(x) = square root ( x^4 -16x^2) Domain: x^4 -16x^2 ≥ 0 -> x^2 ( x^2 - 16) ≥ 0 -> x^2 ≥ 16 => x <= -4, or x ≥ 4 f'(x) = (2x^3 - 16x) / square root ( x^4 -16x^2) c) Find the slope of the line normal to the graph of f at x = 5 Tangent line slope: f'(5) = (2 * 5^3 - 16 * 5)/ sqrt ( 5^4 - 16* 5^2) = 170/15 = 34/3 normal line slope m = -3/34